# What Do Learning Dynamics Reveal About Generalization in LLM Reasoning?

### v1

@@ Abstract
Despite the remarkable capabilities of modern large language models (LLMs), the mechanisms behind their problem-solving abilities remain elusive. In this work, we aim to better understand how the learning dynamics of LLM finetuning shapes downstream generalization. Our analysis focuses on reasoning tasks, whose problem structure allows us to distinguish between memorization (the exact replication of reasoning steps from the training data) and performance (the correctness of the final solution). We find that a model’s generalization behavior can be effectively characterized by a training metric we call pre-memorization train accuracy: the accuracy of model samples on training queries before they begin to copy the exact reasoning steps from the training set. On the dataset level, this metric is able to reliably predict test accuracy, achieving $R^{2}$ of around or exceeding 0.9 across various models (Llama3 8B and Gemma2 9B), datasets (GSM8k, MATH), and training configurations. On a per-example level, this metric is also indicative of whether individual model predictions are robust to perturbations in the training query. By connecting a model’s learning behavior to its generalization, pre-memorization train accuracy can guide targeted improvements to training strategies. We focus on data curation as an example, and show that prioritizing examples with low pre-memorization accuracy leads to 1.5-2x improvements in data efficiency compared to i.i.d. data scaling, and outperforms other standard data curation techniques. Correspondence to: katiekang@eecs.berkeley.edu Code: https://github.com/katiekang1998/reasoning_generalization
@@ A.1 Selection of Memorization Threshold
We find the threshold $p$ by sweeping across a range of values, calculating the pre-memorization train accuracy across different training runs, and selecting the value which yields the strongest predictor of test accuracy. This calibration process only requires a small number of training runs (e.g. 3) to arrive at a robust value of $p$ which can generalize to new training runs on the same model and finetuning dataset. However, it is important the the training runs used for calibration exhibit some spread over test accuracies, and memorization during training.

### v2

@@ Abstract
Despite the remarkable capabilities of modern large language models (LLMs), the mechanisms behind their problem-solving abilities remain elusive. In this work, we aim to better understand how the learning dynamics of LLM finetuning shapes downstream generalization. Our analysis focuses on reasoning tasks, whose problem structure allows us to distinguish between memorization (the exact replication of reasoning steps from the training data) and performance (the correctness of the final solution). We find that a model’s generalization behavior can be effectively characterized by a training metric we call pre-memorization train accuracy: the accuracy of model samples on training queries before they begin to copy the exact reasoning steps from the training set. On the dataset level, this metric is able to reliably predict test accuracy, achieving $R^{2}$ of around or exceeding 0.9 across various models (Llama3 8B, Gemma2 9B), datasets (GSM8k, MATH), and training configurations. On a per-example level, this metric is also indicative of whether individual model predictions are robust to perturbations in the training query. By connecting a model’s learning behavior to its generalization, pre-memorization train accuracy can guide targeted improvements to training strategies. We focus on data curation as an example, and show that prioritizing examples with low pre-memorization accuracy leads to 1.5-2x improvements in data efficiency compared to i.i.d. data scaling, and outperforms other standard data curation techniques. Correspondence to: katiekang@eecs.berkeley.edu Code: https://github.com/katiekang1998/reasoning_generalization
@@ A.1 Selection of Memorization Threshold
We find the threshold $p$ by sweeping across a range of values, calculating the pre-memorization train accuracy across different training runs, and selecting the value which yields the strongest predictor of test accuracy. In Fig. 9, we illustrate how the value of $p$ influences the $R^{2}$ for predicting average test accuracy. We can see that $R^{2}$ degrades smoothly with respect to $p$, which makes it is relatively easy to find a good value of p by sweeping a range of values.
Figure 9: Relationship between the value of p and the coefficient of determination ($R^{2}$) with respect to pre-memorization train accuracy and test accuracy. The $R^{2}$ is taken in aggregate of all the corresponding training runs in Fig. 4.
This calibration process only requires a small number of training runs (e.g. 1-3) to arrive at a robust value of $p$ which can generalize to new training runs on the same model and finetuning dataset, illustrated in Fig. 10. However, it is important the the training runs used for calibration exhibit some spread over test accuracies, and memorization during training.
Figure 10: Calibrating $p$ on a subset of training runs, and evaluating $R^{2}$ on heldout training runs using GSM8k and Llama3 8B. We can see that calibrating on just 1-3 training runs was able to yield a robust value of $p$ which leads to high $R^{2}$ on heldout training runs.
Finally, we also show that the calibration process generalizes to new test examples. We divide the test set into two halves: a calibration test set, and a heldout test set. We calibrate $p$ on the calibration test set, and evaluate the coefficient of determination of the heldout test set. In Fig. 11, we can see that the value of $p$ is able to generalize robustly to new examples on which it had not been calibrated, achieving high coefficient of determination.
Figure 11: Calibrating $p$ using a subset of of the test set (calibration test set), and evaluating $R^{2}$ on a heldout test set using GSM8k and Llama3 8B. We can see that calibrating on just the calibration test set was able to yield a robust value of $p$ which leads to high $R^{2}$ on the heldout test set.